OpenStudy (anonymous):

ind all ordered pairs (a,b) of positive integers such that |3^a − 2^b| = 1.

OpenStudy (anonymous):

i can eyeball some pairs, but thats about it lol >.<

myininaya (myininaya):

(1,1) is one I see

OpenStudy (anonymous):

(1,2)

OpenStudy (anonymous):

(2,3).....i dont see anymore nearby....

OpenStudy (anonymous):

(1,1), (1,2), (2,3)

myininaya (myininaya):

i don't see any more the numbers are getting too big

OpenStudy (anonymous):

if there are anymore pairs, a must be even:\[3^a-2^b=1\iff 2^b=3^a-1=(3-1)(3^{a-1}+3^{a-2}+\cdots+3+1)\]\[2^b=2\cdot(3^{a-1}+3^{a-2}+\cdots+3+1)\iff 2^{b-1}=3^{a-1}+3^{a-2}+\cdots+3+1\]

OpenStudy (anonymous):

i dont think there are anymore pairs...

OpenStudy (anonymous):

i think i can show that a would have to be even, and a would have to be odd., which is impossible.

OpenStudy (anonymous):

We will use mathematical induction on n to prove the slightly stronger result that, for each n greater than or equal to 3, n! can be represented as the sum of n distinct divisors of itself, with 1 as one of the divisors.. We have the base case, for n = 3: 3! = 1 + 2 + 3, with 1 as one of the divisors. For the induction step, we assume the inductive hypothesis, that we have a representation for n = k, with 1 as one of the divisors. That is, k! = 1 + d2 + d3 + ... + dk, with 1 < d2 < ... < dk. Then (k + 1)! = (k + 1) + (k + 1)d2 + ... + (k + 1)dk = 1 + k + (k + 1)d2 + ... + (k + 1)dk That is, the sum of k + 1 distinct divisors of (k + 1)!, with 1 as one of the divisors. It follows by induction that, for each n greater than or equal to 3, n! can be represented as the sum of n distinct divisors of itself.

OpenStudy (anonymous):

So we have:\[3^a-2^b=1\iff 3\cdot (3^{a-1})-2\cdot (2^{b-1})=1\]using Bezout's Identity, we also have:\[3x-2y=1\] for some x and y. All the integer solutions to Bezouts Identity are in the form:\[x=1+2k,y=-1-3k\] where k is an integer because:\[3(1+2k)+2(-1-3k)=3+6k-2-6k=1\]so now we have:\[3(1+2k)-2(1+3k)=1\]which tells us \[3^{a-1}=1+2k\] for some k and some a. Looking at this equation mod 2 gives:\[3^{a-1}\equiv 1 \mod 2\]which says a-1 must be even, but above i showed that a was even. which is a contradiction?